# BannerAgency: Advertising Banner Design with Multimodal LLM Agents

### v1

@@ Abstract
Figure 1: BannerAgency supports designs of various sizes including three common advertisement sizes: skyscraper ($160\times 600$), leaderboard ($728\times 90$), medium rectangle ($300\times 250$). It supports direct SVG generation or rendering in Figma. Both enable continuing editing at element level. Figures best viewed in Adobe Acrobat.
footnotetext: Corresponding author: Heng Wang (heng.wang@sony.com)
@@ 1 Introduction
Advertising banners constitute an instrumental medium in digital marketing campaigns. These graphical displays efficiently capture viewer attention, communicate strategic messaging, and enhance brand visibility across competitive market environments [55, 37, 10, 60]. Unlike generic images, banner ad images are visually harmonized compositions of multiple creative assets including backdrops, brand-related logos or product images, click-to-action (CTA) buttons, campaign texts, and other decorative elements such as shapes and patterns [49, 61]. Campaign texts further require typography design choices, contributing to a vast search space that demands significant time and effort [16, 63]. In addition, advertisers have various size requirements for different display needs [5] and diverse design requirements tailored to specific target audiences and purposes [15]. As design remains inherently creative and subjective, ensuring human interaction and editing capability is crucial for practical applications [21, 33].
Current models accelerate this banner design process by automating separate aspects: product background generation [63, 68, 19, 16, 12], template-based layout matching [69, 18, 8, 3, 39], layout generation  [25, 61, 73, 49], and automatic typography design [14, 28, 54]. Advanced scene-text rendering models [11, 45, 59, 74] and text-to-image (T2I) models such as DALL-E3 [6] are suboptimal for banner design due to their pixel-based generation paradigm, which produces outputs that are difficult to edit and often contains incoherent text as in  Fig. 2(a). Although TypeR [58] addresses the “typos” from T2I models through detection and correction of incorrectly rendered text, it still constrains the design process to a pixel-based approach, limiting subsequent editability. Recent works [33, 31, 64] explore automatic poster generation from prompts, but banner design requires handling both structured requests and unstructured brand identity like logos. Their data dependency and fine-tuning further limit adaptability to diverse banner sizes and requests. Although these works allow text modification through typography rendering, they merge decorative elements with images, restricting editability (see Fig. 2(a), where COLE-based[33, 31] output’s black rectangle cannot be removed).
@@
Figure 2: Algorithm comparison and BannerAgency’s editability.
@@
To bridge these gaps and deliver a fully automatic yet editable banner design solution, we introduce BannerAgency, a training-free framework powered by multimodal large language models (MLLMs) [52, 4, 44]. Our fundamental idea is to utilize MLLMs as AI agents [17, 67] that simulate how a human design team works collaboratively from conception to implementation [15], leveraging the emergent design knowledge inherited from web-scale training data [41, 27, 65]. With tool calling and memory sharing capabilities of LLM agents [22, 47, 26], we effectively navigate the extensive design search space through our specialized agent architecture. The system comprises: (1) a Strategist that constrains design possibilities by establishing brand guidelines and campaign objectives; (2) a Background Designer responsible for visual backdrop generation; (3) a Foreground Designer that creates blueprints for banner assets with engaging content and harmonized styling and positioning; and (4) a Developer that transforms these specifications into editable components via Figma plugin code or SVG generation (see  Fig. 1 for visual examples of the final designs). Unlike pixel-based methods that produce rastic outputs, BannerAgency generates designs directly within professional tools, enabling seamless editing of major design components—text, typography, buttons, shapes, and layouts—within a familiar interface used by professional designers as illustrated in  Fig. 2(b).
@@ 2.1 Graphic Design Generation
Current approaches to graphic design generation face significant limitations. Pixel-based text-to-image models [6, 7, 56, 72] require cumbersome editing, while layout generation research [76, 53, 38, 43, 29, 30, 57, 13, 42, 32, 66] has not adequately addressed simultaneous element generation and arrangement [35]. Existing solutions in prompt-based poster generation [33, 31, 64] rely on domain-specific training, accept only unimodal text input (not multimodal inputs like logos), and lack flexibility. While LayoutPrompt [41] offers a training-free approach, it addresses only layout without end-to-end capabilities. We propose a training-free agent framework that leverages MLLMs to simultaneously generate and arrange design elements in editable code representation from multimodal inputs, enabling flexible banner creation without domain-specific fine-tuning.
@@ 2.2 Multimodal LLM Agents
MLLMs [44, 77, 1, 40] have recently spurred an emerging wave of multimodal agent research [17, 67, 24]. These agents—autonomous entities that leverage tool-calling capabilities, planning, memory, and goal-directed reasoning—have been applied to various design domains including presentation slides generation [75, 20], image manipulation [65, 70], infographic design [27, 9], and sketch creation [62]. Our work extends this evolving frontier by leveraging MLLMs’ emergent capabilities to pioneer a comprehensive banner design solution that generates high-quality designs adaptable to diverse banner requests while ensuring strong editability through its component-based approach.
@@ 3.2 Background Designer
The Background Designer generates prompts for text-to-image (T2I) tools to create the visual canvas for banner ads. Implemented as a ReAct agent [71], it harmonizes advertiser requests, logo characteristics, and campaign objectives stored in memory to produce appropriate background visuals. The agent utilizes three specialized tools: FindImagePath to check for existing background images provided by advertisers, T2I to generate new visuals, and TextChecker to check if the generated image contains text. For new image generation, the agent employs a self-refinement loop [48, 70] to ensure text-free backgrounds, as text elements in the background layer would complicate the Foreground Designer’s work and potentially introduce difficult-to-remove gibberish text rendered directly into pixels. When text is detected by the TextChecker tool, the agent revises its prompt to discourage text generation and repeats the process for up to five iterations, as illustrated in Fig. 4 where a “Black Friday” background was successfully made text-free. During image generation, the agent selects the closest supported aspect ratio from the limited set available in the T2I tool that exceeds the target dimensions—ensuring higher initial resolution—and subsequently downscales to the target banner dimensions. This approach preserves visual fidelity while ensuring exact sizing for various display contexts. The resulting text-free background serves as an appropriate visual foundation, reserving space for the Foreground Designer to add text, CTA buttons, logos, and decorative elements in subsequent stages.
@@ 3.3 Foreground Designer
The evolution of designs through iterative external feedback and refinement is a staple procedure in professional design [2]. Simulating this practice [48, 70], our memory-augmented iterative design refinement process, illustrated in  Fig. 6, enhances banner ad designs by introducing an external critic Design Reviewer. At initial creation ($t=0$), the Foreground Designer generates the first blueprint based solely on the creative brief (banner request, background image, logo). The Developer renders this blueprint, and the Design Reviewer provides initial feedback, which is summarized and stored in memory. For the first refinement ($t=1$), the Foreground Designer refines the blueprint based on this feedback, comparing the new blueprint with the previous one and summarizing into a modification list (i.e., “Compare & Conclude”). For subsequent refinements ($t>1$), both designer and reviewer access the memory that is continuously augmented with new feedback, renderings, and modification summaries after each iteration, enabling increasingly informed decisions. This memory-augmented approach, where each iteration enriches the shared memory with new design artifacts and critiques, ensures both designer and reviewer maintain awareness of design evolution while continuously aligning with brand objectives. The process concludes when iterations are exceeded or the design is deemed production-ready by the reviewer, resulting in the final refined banner ad image.
@@ 3.4 Developer
The Developer plays a crucial role in converting the Foreground Designer’s blueprint into a banner ad image. We propose two implementations for the Developer: generating SVG code or Figma plugin code. Examples of codes in two formats are included in supplementary.
@@
Figure 7: Same campaign for different groups of audience.
@@ 4 BannerRequest400 Benchmark
(a) Banner quality metrics (GPT-4o, human-validated in Tab. 2).                (b) System preference rates (evaluated by both GPT-4o and humans).
Table 1: Quantitative comparison on BannerRequest400 benchmark. * denotes methods requiring explicit banner descriptions.
Existing design datasets, such as DESIGNERINTENTION [31], are limited to uni-modal designer intentions in text format and lack specific addressing of banner advertisement requirements. To address this gap, we introduce BannerRequest400, the first multimodal benchmark specifically designed for evaluating advertising banner generation systems. This novel benchmark pairs brand logos (visual modality) with corresponding banner design requests (textual modality). We curated 100 advertising design intentions from DESIGNERINTENTION and synthesized corresponding vector logos using Claude3.5 Sonnet [4], followed by expert review and refinement to visual fidelity, avoid bias, and maintain authentic logo aesthetics. Our empirical analysis, illustrated in Fig. 7, demonstrates how identical dental clinic promotion requests produce substantially different designs based on target audience—with professional-oriented banners emphasizing efficiency and results, while child-focused versions feature playful elements and incentives. Leveraging this insight, we employed GPT-4o[52] to systematically expand each intention into four distinct banner requests, each targeting a unique audience-purpose combination, resulting in 400 diverse specifications. To ensure industry relevance, we further extended these 400 requests across 13 standard banner dimensions, yielding a comprehensive evaluation set of 5,200 multimodal banner specifications. For benchmarking text-to-image generation methods, we further enhanced each request with detailed specifications that incorporate logo-specific visual characteristics using GPT-4o. BannerRequest400 enables the first rigorous evaluation of banner generation approaches across diverse design requests, facilitating advancement in automatic advertising banner design generation.
5 Experiments
5.1 Experimental Setup
Algorithms.
For baseline approaches on our BannerRequest400 benchmark, we compare with two powerful pixel-based image generation models including closed-sourced DALL-E3 [6] and open-sourced FLUX.1-schnell [7] which is also our backbone for the T2I tool. We also compare with OpenCOLE [33, 31], a prompt-based graphic design model that relies on fine-tuning. For the backbone of our proposed training-free BannerAgency, we explore two multimodal LLM variants including GPT-4o (2024-10-21) [52] and Claude3.5-Sonnet (2024-10-22) [4] with temperature set as 0.3. We provide more details including all agent profile definitions and tool definitions in the supplementary material.
Metrics.
To evaluate banner design quality, we developed six metrics ( Tab. 2) from extensive literature survey on banner design principles [34, 50, 36, 37, 46, 55]. These metrics capture advertising-specific aspects overlooked by general design metrics [31, 23]: audience alignment, brand integration, call-to-action effectiveness, message clarity, brand consistency, and visual appeal—addressing both marketing functionality and aesthetic requirements critical for digital advertising. For completeness and comparison with prior metrics, we also include results using general design metrics in the supplementary material.
Human study.
We conducted three human evaluations to assess different aspects of BannerAgency. First, a system preference study with 20 participants comparing banner designs generated by five BannerAgency variants across 20 randomly selected requests through pairwise comparisons. Second, a refinement effectiveness evaluation where 15 participants assessed quality progression across four iterations of 20 different banner designs, rating both initial and best versions on a 5-point scale. Third, a human-LLM correlation study validating the alignment between GPT-4o’s automated scoring and human perception across six proposed metrics, with 17-19 participants independently grading 25 representative images per metric. See appendix for details.
Figure 8: Visual comparison on BannerRequest400 benchmark. * denotes methods requiring explicit banner descriptions as input (versus abstract requests) due to limited reasoning capabilities, whose detailed prompts are provided in the supplementary.
@@ 5.1 Experimental Setup
(a) Cross-genre aesthetic adaptation.
(b) Cross-cultural linguistic adaptation.
Figure 10: BannerAgency’s design versatility. We demonstrate the high versatility through (a) cross-genre aesthetic adaptation, and (b) cultural adaptation across diverse market contexts, all without requiring additional training or fine-tuning.
Metric Criteria Pearson Spearman ICC
Table 2: Banner design evaluation metrics and human-LLM correlation. Each metric (rated 1-5) assesses a distinct aspect of banner design quality. Pearson, Spearman, and ICC correlation coefficients between human and GPT-4o ratings show strong agreement (all p $<$ 0.001). TAA: Target Audience Alignment; LPS: Logo Placement Score; CTAE: CTA Effectiveness; CPYQ: Copywriting Quality; BIS: Brand Identity Score; AQS: Aesthetic Quality Score. Detailed rating criteria are provided in supplementary.

### v2

@@ Abstract
BannerAgency: Advertising Banner Design with Multimodal LLM Agents
Heng Wang, Yotaro Shimose, Shingo Takamatsu Sony Group Corporation {heng.wang, yotaro.shimose, shingo.takamatsu}@sony.com
footnotetext: Corresponding author: Heng Wang (heng.wang@sony.com, hengwang.vision@gmail.com)
@@ 1 Introduction
Figure 1: Algorithm comparison and BannerAgency’s editability in banner ad image generation.
Advertising banners constitute an instrumental medium in digital marketing campaigns. These graphical displays efficiently capture viewer attention, communicate strategic messaging, and enhance brand visibility across competitive market environments Robinson et al. (2007); Landa (2016); Chen et al. (2021); Vaddamanu et al. (2022). Unlike generic images, banner ad images are visually harmonized compositions of multiple creative assets including backdrops, brand-related logos or product images, click-to-action (CTA) buttons, campaign texts, and other decorative elements such as shapes and patterns Maheshwari et al. (2019); Vempati et al. (2020). Campaign texts further require typography design choices, contributing to a vast search space that demands significant time and effort Du et al. (2024); Wang et al. (2022). In addition, advertisers have various size requirements for different display needs Bannerwise (2022) and diverse design requirements tailored to specific target audiences and purposes community (2025). As design remains inherently creative and subjective, ensuring human interaction and editing capability is crucial for practical applications Guo et al. (2021); Jia et al. (2023).
Current models accelerate this banner design process by automating separate aspects: product background generation Wang et al. (2022); Yang et al. (2024a); Fong and See (2024); Du et al. (2024); Chen et al. (2025), template-based layout matching Yang et al. (2016); Express (2025); Canva (2025); Alibaba (2020); Li and Yang (2023), layout generation  Hu et al. (2021); Vempati et al. (2020); Zhang et al. (2017); Maheshwari et al. (2019); Lin et al. (2025), and automatic typography design Choi et al. (2018); Iluz et al. (2023); Park et al. (2024). Advanced scene-text rendering models Chen et al. (2024); Liu et al. (2024); Tuo et al. (2024); Zhao et al. (2024) and text-to-image (T2I) models such as DALL-E3 Betker et al. (2023) are suboptimal for banner design due to their pixel-based generation paradigm, which produces outputs that are difficult to edit and often contains incoherent text as in Fig. 1(a). Although TypeR Shimoda et al. (2024) addresses the "typos" from T2I models through detection and correction of incorrectly rendered text, it still constrains the design process to a pixel-based approach, limiting subsequent editability. Recent works Jia et al. (2023); Inoue et al. (2024); Wang et al. (2024) explore automatic poster generation from prompts, but banner design requires handling both structured requests and unstructured brand identity like logos. Their data dependency and fine-tuning further limit adaptability to diverse banner sizes and requests. Although these works allow text modification through typography rendering, they merge decorative elements with images, restricting editability (see Fig.1(a), where COLE-basedJia et al. (2023); Inoue et al. (2024) output’s black rectangle cannot be removed).
@@
To bridge these gaps and deliver a fully automatic yet editable banner design solution, we introduce BannerAgency, a training-free framework powered by multimodal large language models (MLLMs) OpenAI (2024); Anthropic (2024); Liu et al. (2023). Our fundamental idea is to utilize MLLMs as AI agents Wu et al. (2024); Zhang et al. (2024); Islam et al. (2024); Qian et al. (2024) that simulate how a human design team works collaboratively from conception to implementation community (2025), leveraging the emergent design knowledge inherited from web-scale training data Lin et al. (2023a); Huang et al. (2024); Wang et al. (2025). With tool calling and memory sharing capabilities of LLM agents Gupta and Kembhavi (2023); Lu et al. (2024); Hu et al. (2024); Xu et al. (2024), we effectively navigate the extensive design search space through our specialized agent architecture. The system comprises: (1) a Strategist that constrains design possibilities by establishing brand guidelines and campaign objectives; (2) a Background Designer responsible for visual backdrop generation; (3) a Foreground Designer that creates blueprints for banner assets with engaging content and harmonized styling and positioning; and (4) a Developer that transforms these specifications into editable components via Figma plugin code or SVG generation. Unlike pixel-based methods that produce rastic outputs, BannerAgency generates designs directly within professional tools, enabling seamless editing of major design components—text, typography, buttons, shapes, and layouts—within a familiar interface used by professional designers as illustrated in Fig. 1(b).
@@ 2.1 Graphic Design Generation
Current approaches to graphic design generation face significant limitations. Pixel-based text-to-image models Betker et al. (2023); BlackForestLab (2024); Rombach et al. (2022); Zhang et al. (2023) require cumbersome editing, while layout generation research Zheng et al. (2019); O’Donovan et al. (2014); Lee et al. (2020); Lin et al. (2024a); Inoue et al. (2023b, a); Shabani et al. (2024); Cheng et al. (2024); Lin et al. (2023b); Iwai et al. (2024); Weng et al. (2024); Yun et al. (2025) has not adequately addressed simultaneous element generation and arrangement Kikuchi et al. (2024). Existing solutions in prompt-based poster generation Jia et al. (2023); Inoue et al. (2024); Wang et al. (2024) rely on domain-specific training, accept only unimodal text input (not multimodal inputs like logos), and lack flexibility. While LayoutPrompt Lin et al. (2023a) offers a training-free approach, it addresses only layout without end-to-end capabilities. We propose a training-free agent framework that leverages MLLMs to simultaneously generate and arrange design elements in editable code representation from multimodal inputs, enabling flexible banner creation without domain-specific fine-tuning.
@@ 2.2 Multimodal LLM Agents
MLLMs Liu et al. (2023); Zhu et al. (2024); Abdin et al. (2024); Li et al. (2023) have recently spurred an emerging wave of multimodal agent research Durante et al. (2024); Xie et al. (2024); He et al. (2024); Lin et al. (2024b); Li et al. (2024); Koh et al. (2024). These agents—autonomous entities that leverage tool-calling capabilities, planning, memory, and goal-directed reasoning—have been applied to various design domains including presentation slides generation Zheng et al. (2025); Ge et al. (2025), image manipulation Wang et al. (2025); Yang et al. (2024b), infographic design Huang et al. (2024); Chen and Huang (2024), and sketch creation Vinker et al. (2024). Our work extends this evolving frontier by leveraging MLLMs’ emergent capabilities to pioneer a comprehensive banner design solution that generates high-quality designs adaptable to diverse banner requests while ensuring strong editability through its component-based approach.
@@ 3.2 Background Designer
The Background Designer generates prompts for text-to-image (T2I) tools to create the visual canvas for banner ads. Implemented as a ReAct agent Yao et al. (2023), it harmonizes advertiser requests, logo characteristics, and campaign objectives stored in memory to produce appropriate background visuals. The agent utilizes three specialized tools: FindImagePath to check for existing background images provided by advertisers, T2I to generate new visuals, and TextChecker to check if the generated image contains text. For new image generation, the agent employs a self-refinement loop Madaan et al. (2023); Yang et al. (2024b) to ensure text-free backgrounds, as text elements in the background layer would complicate the Foreground Designer’s work and potentially introduce difficult-to-remove gibberish text rendered directly into pixels. When text is detected by the TextChecker tool, the agent revises its prompt to discourage text generation and repeats the process for up to five iterations, as illustrated in Fig.3 where a "Black Friday" background was successfully made text-free. During image generation, the agent selects the closest supported aspect ratio from the limited set available in the T2I tool that exceeds the target dimensions—ensuring higher initial resolution—and subsequently downscales to the target banner dimensions. This approach preserves visual fidelity while ensuring exact sizing for various display contexts. The resulting text-free background serves as an appropriate visual foundation, reserving space for the Foreground Designer to add text, CTA buttons, logos, and decorative elements in subsequent stages.
@@ 3.3 Foreground Designer
The evolution of designs through iterative external feedback and refinement is a staple procedure in professional design ADCreative.ai (2024). Simulating this practice Madaan et al. (2023); Yang et al. (2024b), our memory-augmented iterative design refinement process, illustrated in Fig. 5, enhances banner ad designs by introducing an external critic Design Reviewer. At initial creation ($t=0$), the Foreground Designer generates the first blueprint based solely on the creative brief (banner request, background image, logo). The Developer renders this blueprint, and the Design Reviewer provides initial feedback, which is summarized and stored in memory. For the first refinement ($t=1$), the Foreground Designer refines the blueprint based on this feedback, comparing the new blueprint with the previous one and summarizing into a modification list (ie, "Compare & Conclude”). For subsequent refinements ($t>1$), both designer and reviewer access the memory that is continuously augmented with new feedback, renderings, and modification summaries after each iteration, enabling increasingly informed decisions. This memory-augmented approach, where each iteration enriches the shared memory with new design artifacts and critiques, ensures both designer and reviewer maintain awareness of design evolution while continuously aligning with brand objectives. The process concludes when iterations are exceeded or the design is deemed production-ready by the reviewer, resulting in the final refined banner ad image.
@@ 3.4 Developer
The Developer plays a crucial role in converting the Foreground Designer’s blueprint into a banner ad image. We propose two implementations for the Developer: generating SVG code or Figma plugin code. Examples of codes in two formats are included in Appendix Fig. 10.
@@
Figure 6: Same campaign for different audiences.
@@ 4 BannerRequest400 Benchmark
Existing design datasets, such as DESIGNERINTENTION Inoue et al. (2024), are limited to uni-modal designer intentions in text format and lack specific addressing of banner advertisement requirements. To address this gap, we introduce BannerRequest400, the first multimodal benchmark specifically designed for evaluating advertising banner generation systems. This novel benchmark pairs brand logos (visual modality) with corresponding banner design requests (textual modality). We curated 100 advertising design intentions from DESIGNERINTENTION and synthesized corresponding vector logos using Claude3.5 Sonnet Anthropic (2024), followed by expert review and refinement to visual fidelity, avoid bias, and maintain authentic logo aesthetics. Our empirical analysis, illustrated in Fig.6, demonstrates how identical dental clinic promotion requests produce substantially different designs based on target audience—with professional-oriented banners emphasizing efficiency and results, while child-focused versions feature playful elements and incentives. Leveraging this insight, we employed GPT-4oOpenAI (2024) to systematically expand each intention into four distinct banner requests, each targeting a unique audience-purpose combination, resulting in 400 diverse specifications. To ensure industry relevance, we further extended these 400 requests across 13 standard banner dimensions, yielding a comprehensive evaluation set of 5,200 multimodal banner specifications. For benchmarking text-to-image generation methods, we further enhanced each request with detailed specifications that incorporate logo-specific visual characteristics using GPT-4o. BannerRequest400 enables the first rigorous evaluation of banner generation approaches across diverse design requests, facilitating advancement in automatic advertising banner design generation.
5 Experiments
5.1 Experimental Setup
Algorithms.
For baseline approaches on our BannerRequest400 benchmark, we compare with two powerful pixel-based image generation models including closed-sourced DALL-E3 Betker et al. (2023) and open-sourced FLUX.1-schnell BlackForestLab (2024) which is also our backbone for the T2I tool. We also compare with OpenCOLE Jia et al. (2023); Inoue et al. (2024), a prompt-based graphic design model that relies on fine-tuning. For the backbone of our proposed training-free BannerAgency, we explore two multimodal LLM variants including GPT-4o (2024-10-21) OpenAI (2024) and Claude3.5-Sonnet (2024-10-22) Anthropic (2024) with temperature set as 0.3. We provide more details including all agent profile definitions and tool definitions in Appendix A.1.
Metrics.
To evaluate banner design quality, we developed six metrics (Tab. 1) from extensive literature survey on banner design principles Katai (2024); Malhotra (2024); King (2020); Landa (2016); Lohtia et al. (2003); Robinson et al. (2007). These metrics capture advertising-specific aspects overlooked by general design metrics Inoue et al. (2024); Haraguchi et al. (2024): audience alignment, brand integration, call-to-action effectiveness, message clarity, brand consistency, and visual appeal—addressing both marketing functionality and aesthetic requirements critical for digital advertising. For completeness and comparison with prior metrics, we also include results using general design metrics in Appendix A.4.
Metric Criteria Pearson Spearman ICC
Table 1: Banner design evaluation metrics and human-LLM correlation. Each metric (rated 1-5) assesses a distinct aspect of banner design quality. Pearson, Spearman, and ICC correlation coefficients between human and GPT-4o ratings show strong agreement (all p << 0.001). TAA: Target Audience Alignment; LPS: Logo Placement Score; CTAE: CTA Effectiveness; CPYQ: Copywriting Quality; BIS: Brand Identity Score; AQS: Aesthetic Quality Score. Detailed rating criteria are provided in supplementary.
Human study.
We conducted three human evaluations to assess different aspects of BannerAgency. First, a system preference study with 20 participants comparing banner designs generated by five BannerAgency variants across 20 randomly selected requests through pairwise comparisons. Second, a refinement effectiveness evaluation where 15 participants assessed quality progression across four iterations of 20 different banner designs, rating both initial and best versions on a 5-point scale. Third, a human-LLM correlation study validating the alignment between GPT-4o’s automated scoring and human perception across six proposed metrics, with 17-19 participants independently grading 25 representative images per metric. See Appendix A.2 for details.
@@ 5.1 Experimental Setup
(a) Banner quality metrics (GPT-4o, human-validated in Tab. 1).       (b) System preference rates by GPT-4o and humans.
Table 2: Quantitative comparison on BannerRequest400 benchmark.
Figure 7: Visual results. * denotes methods requiring explicit banner descriptions, whose detailed prompts are provided in Appendix A.5. More visual results can be found in Appendix A.7.
@@
Figure 9: BannerAgency’s design versatility. We demonstrate the high versatility through cross-genre aesthetic adaptation (top), and cultural adaptation (bottom) across diverse market contexts, all without requiring additional training or fine-tuning.